OpenStudy (anonymous):

Which equation is a quadratic equation? y = 2x(5 − x) + 7x y = 8x + 2 y + x2 = (x − 2)(x + 8) y − 7x = (x2 + 1)(x − 3)

OpenStudy (anonymous):

which one has only x^2 terms and below

OpenStudy (anonymous):

C?

OpenStudy (anonymous):

actually this question is a bit tricky since the third one looks quadratic, but actually is not

OpenStudy (anonymous):

\[y + x^2 = (x − 2)(x + 8) \] \[y+x^2=x^2+6x-16\] and the \(x^2\) term cancels from both sides to give \[y=6x-16\] so it is not quadratic

OpenStudy (anonymous):

:/

OpenStudy (anonymous):

multiply out for the first one, combine like terms, and you will see that it is a quadratic

OpenStudy (anonymous):

oh yea, duh! I should've seen that. Thank you

OpenStudy (anonymous):

yw